OpenStudy (anonymous):

Solve by the elimination method. (Please show your work so I may understand how you got the answer. WILL FAN AND MEDAL :) 9x + 2y = 2 21x + 6y = 4

OpenStudy (kl0723):

you will need to multiply your first equation by -3: -27x +-6y = -6 21x + 6y = 4 cancel -6+6=0 combine the equations: -6x=2 x=2/-6 x=-1/3 now plug in this value of x into one of the main equations: 9(-1/3)+2y=2 2y=2+3 y=5/2

OpenStudy (anonymous):

Thank You Guys :)!

OpenStudy (kl0723):

wait... is actually 1/3

OpenStudy (anonymous):

How ?

OpenStudy (studygurl14):

times the first equation by -3 to get -27x - 6y = -6 Then add the other equation to get -6x + 0y = -2 --> -6x = -2 divide x = 1/3 Plug in 9(1/3) + 2y = 2 3 + 2y = 2 2y = -1 y = -1/2 (1/3, -1/2)

OpenStudy (kl0723):

that's right :) @StudyGurl14 made a mistake with the signs

OpenStudy (studygurl14):

@kl0723 i didn't make the mistake. you did. just sayin

OpenStudy (anonymous):

Oh Okay... Thank Guys Your Awesome! ^_^ Kudos to you both!

OpenStudy (kl0723):

@StudyGurl14 that's what I meant. just saying

OpenStudy (studygurl14):

lol sorry i misintrepetted @kl0723

OpenStudy (kl0723):

:P

OpenStudy (anonymous):

Lol play nice guys and once again KUdos!